### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable success and demonstrated superior performance across various tasks, including natural language processing (NLP), weather forecasting, biological protein folding, text generation, and solving mathematical problems. However, many real-world data exhibit highly non-Euclidean latent hierarchical anatomy, such as protein networks, transportation networks, financial networks, brain networks, and linguistic structures or syntactic trees in natural languages. Effectively learning intrinsic semantic entailment and hierarchical relationships from these raw, unstructured input data using LLMs remains an underexplored area. Due to its effectiveness in modeling tree-like hierarchical structures, hyperbolic geometry – a non-Euclidean space – has rapidly gained popularity as an expressive latent representation space for complex data modeling across domains such as graphs, images, languages, and multi-modal data. Here, we provide a comprehensive and contextual exposition of recent advancements in LLMs that leverage hyperbolic geometry as a representation space to enhance semantic representation learning and multi-scale reasoning. Specifically, the paper presents a taxonomy of the principal techniques of Hyperbolic LLMs (HypLLMs) in terms of four main categories: (1) hyperbolic LLMs through exp/log maps; (2) hyperbolic fine-tuned models; (3) fully hyperbolic LLMs, and (4) hyperbolic state-space models. We also explore crucial potential applications and outline future research directions. A repository of key papers, models, datasets, and code implementations is available at  https://github.com/sarangp2402/Hyperbolic-LLM-Models/tree/main.
@@ 1 Introduction
The emergence of foundation models and large language models (LLMs) has significantly empowered modern artificial intelligence (AI) systems, producing substantial impacts across a wide range of application fields, such as healthcare [52, 1], finance [64], computer programming [14, 110], earth system modeling [7, 99], autonomous systems [113], etc. Prevalent LLMs, such as GPT [2], Llama [33], Gemini [98], Claude [4], BERT [53], built on transformer architectures [105], excel at capturing long-range dependencies and perform effectively in various natural language processing (NLP) tasks, such as summarization, answering questions, translation, and generation of coherent and contextually relevant text. Typically, LLMs were designed to learn token embeddings in flat Euclidean space $\mathbb{R}^{n}$, where each token is represented as a single vector, leveraging the attention mechanism to capture rich contextual information. However, Euclidean embeddings present two major limitations: 1) flat Euclidean space (with curvature $\mathcal{K}=0$) has limited capacity to embed large and complex taxonomies with high distortions; 2) they struggle to capture logical and semantic entailment information during representation learning, including latent hierarchical structural features, which are critical for complex semantic reasoning and question answering tasks. This limitation is particularly evident when processing complex multi-modal data that exhibits tree-like structures with multi-scale features at various levels. Furthermore, numerous real-world datasets inherently exhibit hierarchical tree-like structures in non-Euclidean space [77, 57], such as textual entailment ontology data, linguistic structures, and phylogenetic trees of DNA sequences [8]. However, hyperbolic space ($\mathbb{H}^{n}$) with negative curvature ($\mathcal{K}<0$) is well-suited for representing the aforementioned large-scale hierarchical datasets, offering significant advantages over conventional Euclidean space ($\mathbb{R}^{n}$) for three key properties: 1) Hyperbolic space grows exponentially, aligning naturally with the tree-like expansion of hierarchical data, allowing for the effective preservation and representation of hierarchies with lower distortion. 2) Hyperbolic embeddings excel at preserving both local (leaf-level) and global (root-level) relationships, ensuring a comprehensive representation of hierarchical patterns. 3) Hyperbolic space ($\mathbb{H}^{n}$) offers high representational capacity even in low dimensions, reducing computational overhead.
@@
As such, learning high-quality feature representations in hyperbolic space has gained significant interest across various research fields, including graph machine learning, computer vision, and natural language processing. Notably, in the domain of geometric deep learning (GDL), most complex networks are scale-free, characterized by a power-law distribution of node degrees. Hence, prevalent graph neural network (GNN) architectures have been successfully generalized to their hyperbolic variants, including hyperbolic node embedding with shallow skipgram model [11], hyperbolic GNNs [69], hyperbolic graph convolutional networks (GCNs) [13], and fully hyperbolic GNNs [108]. Hyperbolic GNN variants provide a powerful framework for modeling node hierarchies and embedding large-scale graphs into hyperbolic spaces. These approaches, characterized by fewer parameters and enhanced generalization capabilities, have been increasingly applied to various graph representation learning tasks, achieving great performance improvements over traditional Euclidean embeddings. Key applications include knowledge graph completion [57, 55, 12, 109, 114], and brain network embedding for detecting neurological degenerative diseases and predicting the age of the human brain [5, 90]. Yang et al. [121] provide a comprehensive review of hyperbolic graph neural network methods and their applications across multiple domains. However, despite these advances in hyperbolic embeddings for specific domains, a comprehensive understanding of how hyperbolic geometry can be systematically integrated into large language models remains fragmented across the literature.
@@
In the NLP domain, language data typically exhibits a Zipfian distribution, as described by Zipf’s law [85], which reflects its hierarchical structure and motivates the use of hyperbolic geometry for representation. Petrovski [84] was the first to demonstrate that embedding sentences via simple Möbius averaging in the Poincaré disk outperforms Euclidean averaging on standard textual-entailment benchmarks, showing that negative curvature can benefit natural-language inference even with shallow RNN encoders. To capture useful word-level hierarchies for a variety of NLP tasks, hyperbolic geometry (i.e., Poincaré ball model) has been used to enhance conventional word embedding methods (e.g., word2vec and GloVe) [77, 100]. In particular, hyperbolic word embeddings have demonstrated superior performance compared to their Euclidean counterparts in lexical entailment prediction, achieving greater efficiency and improved generalization performance. Furthermore, hyperbolic word embeddings have demonstrated exceptional performance in word similarity, analogy, and hypernymy detection by using Gaussian mappings [100] in the Riemannian manifold. To extend hyperbolic embedding approaches to “deep” sequence models such as recurrent neural networks (RNNs), [39] first proposed generalizing vanilla RNNs and gated recurrent units (GRUs) to hyperbolic spaces, where the authors leveraged the theory of gyrovector spaces to reformulate Euclidean softmax regression, feedforward layers, RNNs, and GRUs, enabling them to operate on the Poincaré ball. The results demonstrate significant improvements in the performance of natural language inference (e.g., textual entailment) and noisy-prefix recognition. Fully hyperbolic neural networks [16] for NLP tasks represent a significant advancement by operating entirely within hyperbolic space using Lorentz transformations, overcoming the limitations of previous hybrid approaches that relied on tangent space operations.
Moreover, recent empirical investigations have revealed that pre-trained language models exhibit inherent hyperbolic characteristics that strongly support geometric approaches to language representation. For example, token frequency distributions in LLMs follow power-law patterns with exponents around $\alpha\approx 1.9$, where high-frequency abstract tokens cluster near the origin while low-frequency specific terms are positioned farther away, and the neural networks-derived token embeddings demonstrate significant hyperbolicity with hyperbolicity values of $\delta\approx 0.08-0.12$ across multiple LLM architectures [119]. These findings provide compelling empirical evidence that hyperbolic geometry aligns naturally with the intrinsic structure of language model representations, validating the theoretical foundations for hyperbolic LLMs development.
@@
Herein, we primarily focus on recent advances in developing powerful LLMs that leverage hyperbolic geometry as their representation space, aiming to further enhance expressiveness, improve reasoning, capture hierarchical structures, and optimize efficiency for large-scale complex data. While the dominant paradigm has been Transformer-based architectures like GPT-4, Llama-3, Gemini, Claude-3, and BERT [105], which excel as LLMs, the field has also witnessed the emergence of alternative architectures such as state-space models (SSMs), with Mamba being a prominent example that offers improved computational efficiency over long-range dependencies [41]. Mamba’s improved successor Mamba2 [26] has also shown strong performance implementing state-space duality between attention and state space modeling. Both Transformer and Mamba architectures present unique opportunities and challenges when adapted to hyperbolic geometry, which we explore throughout this paper. We use the term LLMs broadly to encompass modern foundation models across diverse architectures and modalities, not limited to language data, but on extensive datasets to achieve strong performance across multiple downstream tasks. In Section 2, we first present the foundations of hyperbolic geometry, including hyperbolicity measures, hyperbolic models, and key optimization techniques in the hyperbolic manifold. Section 3 presents a taxonomy of existing hyperbolic LLMs approaches. Section 4 provides a benchmarking analysis of hyperbolic LLMs, evaluating their performance on mathematical reasoning, hierarchical language reasoning, while identifying current benchmarking limitations and the need for standardized evaluation frameworks. Section 5 explores emerging applications of hyperbolic LLMs across three primary domains: natural language processing, computer vision, and multi-modal visual-semantic representation and inference. Finally, we conclude the paper with a discussion on future research directions.
@@ 2 Foundations of hyperbolic geometry
Hyperbolic space $\mathbb{H}^{n}$ is a non-Euclidean geometric space with constant sectional negative curvature ($\mathcal{K}<0$), in which the parallel postulate fails, i.e., allowing multiple parallel lines to pass through a point outside a given line. Hyperbolic space can be seen as a “continuous version of trees”, characterized by negative curvature that makes distances grow exponentially toward the boundary. Notably, hyperbolic space shares a deep connection with Minkowski spacetime in special relativity, as the Lorentz model of hyperbolic geometry is isometric to the hyperboloid model used in Minkowski space [6, 78]. Figure 1 illustrates the three fundamental geometric spaces that underpin different representation learning approaches. Euclidean space ($\mathcal{K}=0$) serves as the foundation for traditional LLMs due to its computational simplicity and compatibility with standard linear algebra operations. Spherical geometry ($\mathcal{K}>0$) finds applications in constraining embeddings to unit spheres and modeling globally compact topologies. Hyperbolic space ($\mathcal{K}<0$), distinguished by its exponential volume growth, provides the geometric foundation for HypLLMs. Moreover, hyperbolic embeddings can represent hierarchical structures much more parsimoniously – for instance, Nickel and Kiela [77] showed that a 2-dimensional hyperbolic model captured the WordNet noun hierarchy better than a 200-dimensional Euclidean model, highlighting the vastly greater representational capacity of negatively curved space.
@@
Figure 1: Three commonly used representation spaces: Euclidean space, where traditional LLMs operate; spherical geometry, which constrains embeddings to bounded surfaces; and hyperbolic space, the basis of HypLLMs, which naturally models hierarchical structures.
2.1 Hyperbolicity measure
$\delta$-hyperbolicity introduced by Gromov [40] quantifies how closely a space resembles a tree-like structure, where $\delta\geq 0$ measures the deviation from perfect tree-like geometry. For any 4-tuple $(x,y,z,t)$ in a metric space, $\delta(x,y,z,t)$ is defined as half the difference between the biggest two of the following sums:
(1) $d(x,y)+d(z,t),\quad d(x,z)+d(y,t),\quad d(x,t)+d(y,z).$
The space’s $\delta$-hyperbolicity is defined as the supremum of these values across all possible 4-tuples. Smaller $\delta$ values indicate a higher similarity to tree structures, making this hyperbolic space more naturally suited for hierarchical modeling [77, 38].
@@ 2.2 Hyperbolic representation models
Two commonly used isomorphic models for representing hyperbolic space, the Poincaré ball ($\mathcal{B}_{r}^{n}$) model and the Lorentz model (or hyperboloid manifold - $\mathcal{L}_{r}^{n+1}$), are illustrated in the Figure 2. The sectional curvature of hyperbolic space is determined by $\mathcal{K}=-1/r^{2}$, where $r>0$ is a scale parameter (or radius) that determines the amount of curvature, i.e., larger $r$ $\rightarrow$ smaller curvature ($\mathcal{K}$) $\rightarrow$ flatter space; $\mathcal{K}=0$ is Euclidean space when $r\rightarrow\infty$. In the standard hyperbolic space with sectional curvature $\mathcal{K}=-1$, the curvature parameter $r=1$. Both $\mathcal{B}_{r}^{n}$ and $\mathcal{L}_{r}^{n+1}$ represent the same abstract $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ with radius $r$, but use different coordinate systems. The Poincaré ball $\mathcal{B}_{r}^{n}$ provides a bounded $n$-dimensional representation, while the Lorentz model $\mathcal{L}_{r}^{n+1}$ embeds the hyperbolic space as a hyperboloid in $(n+1)$-dimensional Minkowski spacetime.
Figure 2: Illustration of the two primary hyperbolic representation models used in HypLLMs: the Lorentz model (top left) and the Poincaré ball model (right). The Lorentz model represents hyperbolic space on a single sheet of hyperboloid embedded in Minkowski space $\mathcal{L}_{r}^{n+1}$, offering closed-form geodesic computations and improved numerical stability. The Poincaré ball in space $\mathcal{B}_{r}^{n}$ is a conformal model where geodesics are circular arcs and points near the boundary encode fine-grained leaf-level hierarchy. The lower oval shows the tangent space used in log/exp-based manifold operations for optimization and model computation. These geometric representations provide the foundation for hierarchical embeddings in large-scale hyperbolic LLMs.
@@
The Poincaré ball model represents $\mathbb{H}^{n}$ as the open unit ball $\mathcal{B}_{r}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<r\}$ (where $\|\cdot\|_{2}$ denotes the Euclidean norm) with metric $ds^{2}=\lambda_{x}^{2}\,d\mathbf{x}^{2}$, where $d\mathbf{x}^{2}$ is the Euclidean metric and $\lambda_{x}=\frac{2}{1-\|x\|^{2}}$ is the conformal factor that scales distances according to the hyperbolic geometry [77]. Geodesics in this manifold are arcs that tend toward the boundary $\|x\|\to 1$. The geodesic distance between two points $u,v\in\mathbb{B}^{n}$ is:
(2) $d_{\mathbb{B}}(u,v)\;=\;\mathrm{arcosh}\!\Big{(}1+2\,\frac{\|u-v\|^{2}}{(1-\|u\|^{2})(1-\|v\|^{2})}\Big{)},$
@@
Note that the Lorentz model uses only the upper sheet ($x_{0}>0$) of the two-sheeted hyperboloid. This upper sheet constraint ensures that all points represent valid hyperbolic space coordinates. The hyperbolic distance in this model is
(4) $d_{\mathcal{L}}(x,y)\;=\;\mathrm{arcosh}\!\big{(}-\langle x,y\rangle_{L}\big{)},$
and one can convert between the Lorentz and Poincaré representations via explicit coordinate mappings [76]. Both models are useful; notably, the Lorentz model often provides better numerical stability for optimization (since it spreads points more evenly when far from the origin). In either model, gradient-based learning is performed with Riemannian optimization techniques (e.g., RSGD) or projected Euclidean updates via the exponential and logarithmic maps.

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success and demonstrated superior performance across various tasks, including natural language processing (NLP), weather forecasting, biological protein folding, text generation, and solving mathematical problems. However, many real-world data exhibit highly non-Euclidean latent hierarchical anatomy, such as protein networks, transportation networks, financial networks, brain networks, and linguistic structures or syntactic trees in natural languages. Effectively learning intrinsic semantic entailment and hierarchical relationships from these raw, unstructured input data using LLMs remains an underexplored area. Due to its effectiveness in modeling tree-like hierarchical structures, hyperbolic geometry – a non-Euclidean space – has rapidly gained popularity as an expressive latent representation space for complex data modeling across domains such as graphs, images, languages, and multi-modal data. Here, we provide a comprehensive and contextual exposition of recent advancements in LLMs that leverage hyperbolic geometry as a representation space to enhance semantic representation learning and multi-scale reasoning. Specifically, the paper presents a taxonomy of the principal techniques of Hyperbolic LLMs (HypLLMs) in terms of four main categories: (1) hyperbolic LLMs through exp/log maps; (2) hyperbolic fine-tuned models; (3) fully hyperbolic LLMs, and (4) hyperbolic state-space models. We also explore crucial potential applications and outline future research directions. A repository of key papers, models, datasets, and code implementations is available at  https://github.com/sarangp2402/Hyperbolic-LLM-Models.
@@ 1 Introduction
The emergence of foundation models and large language models (LLMs) has significantly empowered modern artificial intelligence (AI) systems, producing substantial impacts across a wide range of application fields, such as healthcare [54, 1], finance [65], computer programming [15, 114], earth system modeling [8, 103], autonomous systems [117], etc. Prevalent LLMs, such as GPT [2], Llama [34], Gemini [102], Claude [4], BERT [55], built on transformer architectures [109] perform effectively in various natural language processing (NLP) tasks, such as summarization, answering questions, translation, and generation of coherent and contextually relevant text. Typically, LLMs were designed to learn token embeddings in flat Euclidean space, where each token is represented as a single vector, leveraging the attention mechanism to capture rich contextual information. However, Euclidean embeddings present two major limitations: 1) flat Euclidean space (with curvature $\mathcal{K}=0$) has limited capacity to embed large and complex taxonomies with high distortions; 2) they struggle to capture logical and semantic entailment information during representation learning, including latent hierarchical structural features, which are critical for complex semantic reasoning and question answering tasks. This limitation is particularly evident when processing complex multi-modal data that exhibits tree-like structures with multi-scale features at various levels. Furthermore, numerous real-world datasets inherently exhibit hierarchical tree-like structures in non-Euclidean space [80, 59], such as textual entailment ontology data, linguistic structures, and phylogenetic trees of DNA sequences [9]. However, hyperbolic space with negative curvature ($\mathcal{K}<0$) is well-suited for representing the aforementioned large-scale hierarchical datasets, offering significant advantages over conventional Euclidean space for three key properties: 1) Hyperbolic space grows exponentially, aligning naturally with the tree-like expansion of hierarchical data, allowing for the effective preservation and representation of hierarchies with lower distortion. 2) Hyperbolic embeddings excel at preserving both local (leaf-level) and global (root-level) relationships, ensuring a comprehensive representation of hierarchical patterns. 3) Hyperbolic space offers high representational capacity even in low dimensions, reducing computational overhead.
@@
Learning high-quality representations in hyperbolic space has gained growing interest in graph machine learning, computer vision, and natural language processing. In geometric deep learning, many real-world networks are scale-free, motivating the development of hyperbolic extensions of traditional GNNs, such as hyperbolic skip-gram models [12], hyperbolic GNNs [71], hyperbolic graph convolutional networks (GCNs) [14], and fully hyperbolic GNNs [112]. These models effectively capture hierarchical structures and large-scale graph geometry with fewer parameters and stronger generalization, showing superior performance in tasks like knowledge graph completion [59, 57, 13, 113, 118], and brain network embedding for disease detection and brain age prediction [5, 94]. Previous works [125, 87] provide a comprehensive review of hyperbolic deep neural networks or hyperbolic graph neural network methods and their applications across multiple domains. However, despite these advances in hyperbolic embeddings for specific domains, a comprehensive understanding of how hyperbolic geometry can be systematically integrated into large language models remains fragmented across the literature.
@@
For natural language data, the word frequency typically follows a Zipfian distribution, as described by Zipf’s law [89], which reflects its hierarchical structure and motivates the use of hyperbolic geometry for representation. Petrovski [88] demonstrated that embedding sentences in the Poincaré disk using Möbius averaging outperforms Euclidean averaging on textual-entailment benchmarks, showing that negative curvature can benefit natural-language inference. To capture useful word-level hierarchies for a variety of NLP tasks, hyperbolic geometry (i.e., Poincaré ball model) has been used to enhance conventional word embedding methods (e.g., word2vec and GloVe) [80, 104]. In particular, hyperbolic word embeddings outperform their Euclidean counterparts in lexical entailment prediction, offering higher efficiency and better generalization. They also achieve exceptional performance in word similarity, analogy, and hypernymy detection using Gaussian mappings on the Riemannian manifold [104]. Several architectures extend RNNs and GRUs into hyperbolic space via gyrovector operations or Lorentz transformations, enabling end-to-end training without returning to tangent space [40, 17]. The results demonstrate significant improvements in the performance of natural language inference and noisy-prefix recognition. Fully hyperbolic neural networks [17] for NLP tasks represent a significant advancement by operating entirely within hyperbolic space using Lorentz transformations, overcoming the limitations of previous hybrid approaches that relied on tangent space operations.
@@
More recently, pre-trained LLMs have been found exhibit intrinsic hyperbolic structure, with token frequencies following a power law ($\alpha\approx 1.9$) and embeddings showing tree-like organization ($\delta$-hyperbolicity $\approx 0.08-0.12$[123]. These findings provide compelling empirical evidence that hyperbolic geometry aligns naturally with the intrinsic structure of language model representations, validating the theoretical foundations for hyperbolic LLMs (HypLLMs) development. Herein, we primarily focus on recent advances in developing powerful LLMs that leverage hyperbolic geometry as their representation space, aiming to further enhance expressiveness, improve reasoning, capture hierarchical structures, and optimize efficiency for large-scale complex data. Note that we use the term “large language models” (LLMs) broadly to encompass neural network architectures, primarily but not exclusively based on transformers, that are trained on extensive datasets to achieve strong performance across different tasks. While initially developed for natural language processing, modern LLMs include foundation models operating on diverse modalities such as vision, audio, and multimodal data. Modern LLMs operate over vocabularies on the order of 50K-100K+ tokens [100], yet the complex semantic hierarchies within the data is not well-captured by Euclidean embeddings, which have reduced representation capacity for hierarchical data compared to hyperbolic alternatives. While the dominant paradigm has been Transformer-based architectures, which excel as LLMs, the field has also witnessed the emergence of alternative architectures such as state-space models (SSMs), with Mamba being a prominent example that offers improved computational efficiency over long-range dependencies [42]. Mamba’s improved successor Mamba2 [27] has also shown strong performance implementing state-space duality between attention and state space modeling. Both Transformer and Mamba architectures present unique opportunities and challenges when adapted to hyperbolic geometry, which we explore throughout this paper. This paper serves as a comprehensive survey that synthesizes and unifies the fragmented landscape of hyperbolic geometry integration in large language models. Our primary goal is to provide a coherent mathematical and architectural framework of existing approaches, clarifying their relationships, and establishing a foundation for future research in this emerging field. In Section 2, we first present the foundations of hyperbolic geometry, representation models, manifold mappings, and key optimization techniques in the hyperbolic manifold. Section 3 presents a taxonomy of existing hyperbolic LLM approaches. Section 4 provides a benchmarking analysis of hyperbolic LLMs, evaluating their performance on mathematical reasoning and hierarchical language reasoning, while identifying current benchmarking limitations and the need for standardized evaluation frameworks. We also discuss the key challenges in building hyperbolic LLMs in this section. Section 6 explores emerging applications of hyperbolic LLMs across three major domains: natural language processing, computer vision, and multi-modal visual-semantic representation and inference. Finally, we conclude the paper with a discussion on future research directions.
@@ 2 Foundations of hyperbolic geometry
Hyperbolic space $\mathbb{H}^{n}$ is a non-Euclidean geometric space with constant sectional negative curvature ($\mathcal{K}<0$), where the parallel postulate fails, i.e., allowing multiple parallel lines to pass through a point outside a given line. It can be viewed as a “continuous version of trees”, characterized by negative curvature that makes distances grow exponentially toward the boundary. Notably, hyperbolic space shares a deep connection with Minkowski spacetime in special relativity, as the Lorentz model of hyperbolic geometry is isometric to the hyperboloid model used in Minkowski space [6, 81]. The three fundamental geometric spaces that underpin different representation learning approaches are Euclidean ($\mathcal{K}=0$, traditional LLMs), spherical ($\mathcal{K}>0$, compact topology), and hyperbolic ($\mathcal{K}<0$, hierarchical data) as represented in Figure 1. Moreover, hyperbolic embeddings can represent hierarchical structures much more parsimoniously – for instance, Nickel and Kiela [80] showed that a 5-dimensional hyperbolic model captured the WordNet noun hierarchy better than a 200-dimensional Euclidean model, highlighting the vastly greater representational capacity of negatively curved space.
@@
Figure 1: Three commonly used representation spaces: Euclidean space, where traditional LLMs operate; spherical geometry, which constrains embeddings to bounded surfaces; and hyperbolic space, the basis of Hyperbolic LLMs, which naturally models hierarchical structures.
@@ 2.1 Hyperbolic representation models
Two commonly used isomorphic models for representing hyperbolic space, the Poincaré ball and the Lorentz model (or hyperboloid manifold), are described in the Figure 2. The sectional curvature of hyperbolic space is determined by $\mathcal{K}=-1/r^{2}$, where $r>0$ is the radius that determines the amount of curvature, i.e., larger $r$ $\rightarrow$ smaller curvature ($\mathcal{K}$) $\rightarrow$ flatter space; $\mathcal{K}=0$ is Euclidean space when $r\rightarrow\infty$. In the standard hyperbolic space with sectional curvature $\mathcal{K}=-1$, the curvature parameter $r=1$. Both $\mathcal{B}_{r}^{n}$ and $\mathcal{L}_{r}^{n+1}$ represent the same abstract $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ with radius $r$, but use different coordinate systems. The Poincaré ball $\mathcal{B}_{r}^{n}$ provides a bounded $n$-dimensional representation, while the Lorentz model $\mathcal{L}_{r}^{n+1}$ embeds the hyperbolic space as a hyperboloid in $(n+1)$-dimensional Minkowski spacetime.
Figure 2: Illustration of the two primary hyperbolic representation models used in HypLLMs: the Lorentz model (top left) and the Poincaré ball model (right). The Lorentz model represents hyperbolic space on a single sheet of hyperboloid embedded in Minkowski space $\mathcal{L}_{r}^{n+1}$, offering closed-form geodesic computations and improved numerical stability. The lower oval shows Poincaré ball disk projection in space $\mathcal{B}_{r}^{n}$, this a conformal model where geodesics are circular arcs and points near the boundary encode fine-grained leaf-level hierarchy. These geometric representations provide the foundation for hierarchical embeddings in hyperbolic LLMs.
@@
An $n$-dimensional Poincaré ball model represents $\mathbb{H}^{n}$ as the open unit ball $\mathcal{B}_{r}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<r\}$ (where $\|\cdot\|$ denotes the Euclidean norm) with metric $ds^{2}=\lambda_{x}^{2}\,dx^{2}$, where $dx^{2}$ is the Euclidean metric and $\lambda_{x}=\frac{2}{1-\|x\|^{2}}$ is the conformal factor that scales distances according to the hyperbolic geometry [80]. Geodesics in this manifold are arcs that tend toward the boundary $\|x\|\to 1$. The geodesic distance between two points $u,v\in\mathcal{B}_{r}^{n}$ is:
(1) $d_{\mathcal{B}}(u,v)=r\cdot\mathrm{arcosh}\!\Big(1+2\,\frac{\|u-v\|^{2}}{(r^{2}-\|u\|^{2})(r^{2}-\|v\|^{2})}\Big),$
@@
Note that the Lorentz model uses only the upper sheet ($x_{0}>0$) of the two-sheeted hyperboloid. This upper sheet constraint ensures that all points represent valid hyperbolic space coordinates. The hyperbolic distance between two points $x$ and $y$ is:
(3) $d_{\mathcal{L}}(x,y)=r\cdot\operatorname{arcosh}\left(\frac{-\langle x,y\rangle_{L}}{r^{2}}\right),$
and one can convert between the Lorentz and Poincaré representations via explicit coordinate mappings [79]. Notably, the Lorentz model often provides better numerical stability for optimization. In either models, gradient-based learning can be performed with Riemannian optimization techniques in Section 2.3 and projected Euclidean updates via the exponential and logarithmic maps.